OpenStudy (anonymous):

What correctly compares 8/15 and 5/8?

OpenStudy (anonymous):

<

OpenStudy (anonymous):

LCM is 120 2^3 * 3 * 5 8/15 = 64/120 5/8 = 75/120 since 75/120 > 64/120, 5/8 > 8/15

OpenStudy (anonymous):

Don't get it. What's the answer?

OpenStudy (anonymous):

Okay I'm gonna go if you're being boring.

OpenStudy (igreen):

What do you mean by compare? Do you mean which one's bigger and which one's smaller?

OpenStudy (igreen):

@OpiGeode

OpenStudy (igreen):

8/15 = 5.33 5/8 = 6.25 I think you can tell which is bigger and which is smaller now :) @OpiGeode

OpenStudy (anonymous):

I didn't write it, this was a test. And I failed it.